OpenStudy (bmxlifestyle):

what is the solution of the system 5x-y=-7 and 3x-y=-2 A -2.5,-5.5 B -2.5,11 C 5, 2.5 D 2.5, 5.5

OpenStudy (bmxlifestyle):

@MikeyMaximum

OpenStudy (anonymous):

x = -5/2 and y = -11/2

OpenStudy (anonymous):

i.e x = -2.5 y = -5.5

OpenStudy (bmxlifestyle):

can u help me with one more